OpenStudy (anonymous):

how to find axis of symmetry? Like for y=x^2-4x-11

OpenStudy (anonymous):

complete the square y=(x-2)^2-7 your axis of symmetry is x=2

OpenStudy (anonymous):

actually its y=(x-2)^2-15 sawwweeee

OpenStudy (anonymous):

Its fine thanks